OpenStudy (anonymous):

MEDAL AND FAN! PLZ HELP Two systems of equations are shown below: System A: 6x+y=2 -x-y=-3 System B: 2x-3y=-10 -x-y=-3 Which of the following statements is correct about the two system equations?

OpenStudy (anonymous):

A) The value of x for System B will be 4 less than the value of x for System A because the coefficient of x in the first equation of System B is 4 less than the coefficient of x in the first equation of System A. B) They will have the same solution because the first equations of both the systems have the same graph. C) They will have the same solution because the first equation of System B is obtained by adding the first equation of System A to 4 times the second equation of System A. D) The value of x for System A will be equal to the value of y for System B because the first equation of System B is obtained by adding -4 to the first equation of System A and the second equations are identical.

OpenStudy (anonymous):

@SolomonZelman I have no idea how to do this can you help me?

OpenStudy (solomonzelman):

I think we first have to solve each of the systems, don't we?

OpenStudy (anonymous):

Alright how do I solve a system :( Could you walk me through one?

OpenStudy (solomonzelman):

Lets start from system A, \(\large\color{black}{ 6x~~~+y=~~~2 \\ -x~-y=~-3}\)

OpenStudy (solomonzelman):

I would suggest to add the equations.

OpenStudy (anonymous):

so like 6x-x+y-y-3=2?

OpenStudy (anonymous):

than it would be 5x-3=2?

OpenStudy (solomonzelman):

\(\large\color{black}{ ~~~~~~6x~~~+y=~~~2 \\^+ \\~~~~~ -x~-y=~-3}\) --------------------- \(\large\color{black}{ ~~~~~~~~~~~~~~~~~~~~~=~~~~~}\)

OpenStudy (anonymous):

add the 3 on both sides?

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

add all of the aXes, all the whYs all the numbers that are on the right sides together.

OpenStudy (anonymous):

so 5x=-1?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So, what is x equal to?

OpenStudy (anonymous):

x=-1/5?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

now that you know that x=-1/5 can you solve for y?

OpenStudy (solomonzelman):

reminding the system is, \(\large\color{black}{ 6x~~~+y=~~~2 \\ -x~-y=~-3}\) ( you can use my latex `\(\large\color{black}{ 6x~~~+y=~~~2 \\ -x~-y=~-3}\)` )

OpenStudy (anonymous):

y=0?

OpenStudy (solomonzelman):

not really, no.

OpenStudy (anonymous):

6y=-1?

OpenStudy (solomonzelman):

sorry, \(\large\color{black}{ 6(-\frac{1}{5})~~~+y=~~~2 \\ -(-\frac{1}{5})~-y=~-3}\) if you get the same result for y, in both of the systems, that will also verify the solution.

OpenStudy (solomonzelman):

\(\large\color{black}{ 6(-\frac{1}{5})~~~+y=~~~2 \\ -(-\frac{1}{5})~-y=~-3}\) \(\large\color{black}{ -6\frac{1}{5}~~~+y=~~~2 \\ \frac{1}{5}~-y=~-3}\)

OpenStudy (anonymous):

double negatives mean positive so 6(-1/5) +(1/5) so it would still be 6(1/5)

OpenStudy (solomonzelman):

wait, I maid a mistake again.

OpenStudy (solomonzelman):

no, I didn't.

OpenStudy (solomonzelman):

sorry for confusing.

OpenStudy (anonymous):

No no its fine I'm here to learn ^_^

OpenStudy (solomonzelman):

\(\large\color{black}{ -6\frac{1}{5}~~~+y \color{red}{+6\frac{1}{5}}=~~~2\color{red}{+6\frac{1}{5}} \\ \frac{1}{5}~-y\color{red}{-\frac{1}{5}}=~-3\color{red}{-\frac{1}{5}}}\)

OpenStudy (solomonzelman):

see what I am doing?

OpenStudy (anonymous):

Yes your adding 6(1/5) to the top numbers and subtracting (1/5) from the bottom numbers

OpenStudy (solomonzelman):

yes, which isolates the y in both of the equations:)

OpenStudy (solomonzelman):

what do you get for y?

OpenStudy (anonymous):

For the top equation I got y=8(1/5) For the bottom equation I got -y=-3(1/5)

OpenStudy (solomonzelman):

yes, I am not sure what do really do here, I went over the work we did: \(\large\color{black}{ 6x~~~+y =~~~2\\ -x~-y=~-3}\) \(\large\color{black}{ 5x=-1}\)

OpenStudy (solomonzelman):

and then x=-1/5.

OpenStudy (solomonzelman):

@Loser66 maybe you see the mistake?

OpenStudy (solomonzelman):

\(\large\color{black}{ 6x+y=2 \\ -x-y=-3}\) \(\large\color{black}{ 6x+y=2 \\ -6x-6y=-18}\) \(\large\color{black}{-5y=-16}\) \(\large\color{black}{ y=5/16}\)

OpenStudy (solomonzelman):